OpenStudy (just_one_last_goodbye):

What is the length of line segment S T if S is at negative six comma negative four and T is at five comma negative nine? Round to the nearest hundredth. ten point seven seven twelve point zero eight thirteen point zero four sixteen point one two

OpenStudy (anonymous):

twelve point zero eight

OpenStudy (just_one_last_goodbye):

u sure?

OpenStudy (anonymous):

yeah . pythagoras sqrt[ (change in x)^2 + (change in y)^2] = sqrt[((-6)-(5))^2 +((-4)-((-9))^2] = sqrt[(-11)^2 + 5^2] = sqrt[121+25] = sqrt[146] since sqrt[144] = 12 and sqrt[146] is slightly more then the answer will be Answer D

OpenStudy (just_one_last_goodbye):

thanks

OpenStudy (anonymous):

no problem :)

OpenStudy (just_one_last_goodbye):

wait why did you put answer D and on the first one it says twelve point zero eight?

OpenStudy (anonymous):

im sorry , that was a mistake. but it is twelve point zero eight .